OpenStudy (anonymous):

help plz!!!!!!!!! A theater offers a 20% discount to senior citizens on all tickets. The regular cost of a matinee ticket is $9.75. Which expression represents the cost of a matinee ticket after the discount is applied? A.9.75 * 1/5 B 9.75 * 4/5 C 9.75 * 6/5 D 9.75 * 9/5

OpenStudy (anonymous):

please help its my first day using openstudy

OpenStudy (anonymous):

i think its b

OpenStudy (anonymous):

\[The~price~of~a~ticket~after~discount=9.75\times \frac{ 100-20 }{ 100 }\]

OpenStudy (anonymous):

are u sure?

OpenStudy (anonymous):

20 percent is the discount price for a ticket

OpenStudy (anonymous):

other people said a

OpenStudy (anonymous):

The question ask to find the cost of a ticket after it is discounted

OpenStudy (anonymous):

\[Discounted~price=\frac{ 20 }{ 100 }\times9.75\]\[=1.95\]

OpenStudy (anonymous):

yea, so A?

OpenStudy (anonymous):

But now the question is asking to find the cost of a ticket after discount.

OpenStudy (anonymous):

Nope. it is not A

OpenStudy (anonymous):

ill just go with b then can u help me with another one

OpenStudy (anonymous):

I will try my best

OpenStudy (anonymous):

During his first year of life, Avery gained 12 3/4 lb. During his second year of life, Avery gained 40% less weight than he did during his first year of life. How much weight did Avery gain during his second year of life? Express your answer as a mixed number in simplest form. lb= btw the answer should be a whole number